### v1

# Language Models Learn Rare Phenomena from Less Rare Phenomena: The Case of the Missing AANNs

Abstract
Language models learn rare syntactic phenomena, but it has been argued that they rely on rote memorization, as opposed to grammatical generalization. Training on a corpus of human-scale in size (100M words), we iteratively trained transformer language models on systematically manipulated corpora and then evaluated their learning of a particular rare grammatical phenomenon: the English Article+Adjective+Numeral+Noun (AANN) construction (“a beautiful five days”). We first compared how well this construction was learned on the default corpus relative to a counterfactual corpus in which the AANN sentences were removed. AANNs were still learned better than systematically perturbed variants of the construction. Using additional counterfactual corpora, we suggest that this learning occurs through generalization from related constructions (e.g., “a few days”). An additional experiment showed that this learning is enhanced when there is more variability in the input. Taken together, our results provide an existence proof that models learn rare grammatical phenomena by generalization from less rare phenomena. Code available at https://github.com/kanishkamisra/aannalysis.
Language Models Learn Rare Phenomena from Less Rare Phenomena:
The Case of the Missing AANNs
Kanishka Misra   Kyle Mahowald Department of Linguistics The University of Texas at Austin {kmisra,kyle}@utexas.edu
Figure 1: A sampling of some of our key experimental manipulations and resulting performance changes. We iteratively train LMs on varied input corpora and measure the learning of the rare aann (“a beautiful five days”) or related constructions. E.g., we train on the default BabyLM corpus, a corpus in which we remove all aanns, a corpus in which we replace all aanns with a corrupted version (“beautiful a five days”) and measure learning of the corrupted version, and a corpus in which we remove aanns and also remove related constructions like “a few weeks is”.
1 Introduction
1.1 Motivation and Prior Work
Humans come to learn and use rare grammatical structures, even if they have encountered those structures only rarely or even not at all (Pullum and Scholz, 2002; Pearl, 2022). For instance, humans accept the grammaticality of the PiPP construction (“surprising though it may be…”) even where the preposed element crosses a finite close boundary (“surprising though I know it may be that…”) (Pullum, 2017) and even though they may plausibly have never encountered such a sentence in their linguistic experience. [^1] Potts (2023) found that the finite-clause-crossing PiPP occurred in only 56 sentences out of a corpus of over 7 billion sentences, which is sufficiently rare that an adult English speaker may never have encountered this construction! How people come to know an utterance is grammatical has occupied a central place in linguistics. Specifically, mastery of never-before-encountered grammatical structures has been taken to mean that people are endowed with innate linguistic knowledge (Chomsky, 1986, 1957, 1965).
Recent evidence, though, suggests that Large Language Models (LLMs) can learn complex grammar (Wilcox et al., 2018; Futrell et al., 2019; Linzen et al., 2016; Mahowald et al., ) even from human-scale amounts of input (Warstadt et al., 2023; Eldan and Li, 2023; Huebner et al., 2021). This raises the possibility that input data, along with an appropriately sophisticated or weakly biased statistical learning mechanism, is sufficient for learning rare constructions by allowing for models to emergently learn appropriate grammatical abstraction (Baroni, 2022; Piantadosi, 2023).
But modern LLMs often have access to much more training input than people do and thus might memorize rare constructions in a way that humans can’t (Warstadt and Bowman, 2022; Warstadt et al., 2023). The possibility that LLMs are “stochastic parrots” (Bender et al., 2021) that rely heavily on memorization, much more so than humans, is a common criticism of using LLMs to study human language (e.g., Chomsky et al., 2023).
There are different levels of memorization, though, requiring different levels of abstraction. Consider the relatively rare AANN construction: “a beautiful five days in Texas” (Solt, 2007; Keenan, 2013; Dalrymple and King, 2019), which is interestingly distinct from the default “five beautiful days in Texas”. A model that strictly memorizes this construction might come to know that “a beautiful five days in Texas” is grammatical but have no idea that “a beautiful four days in Texas” is grammatical if it never appeared in its training. A model that generalizes just a bit more might know that “a beautiful five days in New York” is also grammatical by analogizing that a U.S. state can fill the slot. One that generalizes even more might find “a beautiful three hours at the mall” to be acceptable. Knowing that “an astonishing 200 pages” is acceptable requires generalization beyond mere lexical items. And knowing that “a blue five pencils” is not acceptable (because the adjective is “stubbornly distributive”, Schwarzschild 2011) requires yet more generalization. Even for an idealized learner, it is difficult to precisely formulate how many examples are required to get the right kinds of generalizations for even this one construction.
There is increasing evidence that LLMs can generate novel linguistic utterances (McCoy et al., 2023), and also make subtle judgments on relatively rare constructions like these (Weissweiler et al., 2022; Potts, 2023), including the AANN construction (Mahowald, 2023). If they do so by memorizing examples verbatim from an unrealistically large input corpus, that would not be particularly informative for human processing. But, if they do learn rare grammatical phenomena from smaller amounts of data and can generalize beyond just those verbatim instances, that would raise the question of how they do it. For instance, in the context of the PiPP construction, Potts (2023) speculates that the comparative construction (e.g., “They are happier than we are.”) “may be the key to all of this [i.e., learning the PiPP]” because such constructions are “incredibly common” yet share abstract structure with the PiPP. If LLMs learn rare grammatical structures in part by learning and generalizing structures from much more common constructions, that would be powerful evidence for abstraction in LLMs and raise the possibility that even quite general learners could learn very rare phenomena without strong innate priors.
To that end, our goal in this paper is to study a rare grammatical phenomenon in LMs that are trained on controlled input corpora and that are (a) of human realistic scale, and (b) systematically manipulated with respect to the target constructions as well as key related constructions. Our hypothesis is that generalization abilities of LMs on such rare phenomena come from abstractions and structures learned from more frequent phenomena—and that the more frequent phenomena are the “keys to all of this.”
As a case study, we focus on the aforementioned aann construction, although we highlight how the methods used here could serve as a blueprint for work on other rare grammatical phenomena. Our method is to train different instantiations of a standard transformer model on the 100M-word BabyLM corpus, which we systematically manipulate—via removal and replacement—to shed light on the extent to which frequent and related phenomena encountered during training can give rise to abstractions that allow for generalizations in LMs. To test for generalization, we subjected our different LMs to a series of acceptability tests on sentences which do not appear in the training corpus and which specifically target the special properties of the aann construction. This approach of training on a systematically manipulated corpus has been fruitfully used to debias models (Lu et al., 2020; Maudslay et al., 2019), study whether LMs rely on lexical knowledge in learning syntactic rules (Wei et al., 2021), explore the effect of permuting words on pretrained models (Sinha et al., 2021), and test whether LMs can learn languages judged to be hard for humans (Kallini et al., 2024), among others. Our experiments yielded three main findings, summarized below.
1.2 Summary of findings
First, we find BabyLM-trained LMs to successfully generalize to novel instances of the aann construction. Performance unsurprisingly drops for LMs that were trained without being exposed to even a single aann during training, but perhaps surprisingly, not by all that much—they are substantially above chance. This suggests that certain items present in the training data might give rise to LMs’ non-trivial performance in judging acceptability of aanns. This finding is further strengthened by the fact that LMs trained on counterfactual variants of the aann—e.g., anan and naan, obtained by shuffling word order and are far less likely to share structure with training data items—are unable to generalize to those constructions as well as they do to aanns.
Next, we investigated what might enable LMs’ learning of the aann, by further systematically manipulating their training data to hold out utterances conforming to specific linguistic and statistical phenomena. Through our manipulations, we find LMs become increasingly worse at predicting novel instances of the aann as more frequent, non-AANN-but-AANN-related phenomena are held out. For example, phenomena such as the treatment of measure noun phrases as singular (e.g., a few days is all we need)—similar to how aanns treat a plural NP as singular—end up making unseen aanns less likely by 31% on average. Importantly, these results could not be explained simply by loss of data—LMs that were trained with these phenomena left in but without an equivalently large chunk of the training data removed were almost as good as LMs that never saw aanns. This further strengthens the conclusion that the hypothesized linguistic phenomena did indeed affect generalization of the targeted construction. Notably, LMs are largely affected by these manipulations when they do not see any aanns during training, highlighting the expected non-trivial role of encountering some instances of aanns to show stronger generalization.
Finally, we characterized the aforementioned interplay between the properties of the encountered aanns and the LMs generalizations on novel instances. Here we found LMs that observed aanns with more variability on the adjective, numeral, and noun slots to show better generalization than did LMs that saw more restricted-but-repeating instances of aanns. This importantly mimicked analogous findings of inductive inference in humans across linguistics (Goldberg, 1995, 2005; Suttle and Goldberg, 2011; Baayen, 2009; O’Donnell, 2015) and cognitive psychology studies (Osherson et al., 1990; Xu and Tenenbaum, 2007).
Taken together, these results provide an existence proof that a weakly biased but sophisticated general-purpose statistical learner can in fact learn rare and complex linguistic phenomena, and that this can arise as a result of the presence of key related constructions during training.
2 General Methods
Here we describe methods that we use to characterize learning of the targeted rare construction, the aann. As mentioned in Section 1, we take these to be a general blueprint which can be used to study other rare constructions.
2.1 Data and Model
Corpus
Throughout the paper, we use the BabyLM-strict corpus (Warstadt et al., 2023) as our main training set. We use this primarily because of its tractable size (100M tokens), which allows us to (1) detect and control the instances of the target construction as well as related linguistic phenomena; and (2) train a whoppingly large amount of LMs in sufficient time. Since much of our experiments involve ablating constructions present in this corpus, a majority of the LMs are trained on ablated versions of BabyLM.
Construction Detection
We use regexes over a POS-tagged version of the BabyLM training data to detect aann instances. [^2] We use spaCy for getting POS-tags, see Appendix B for the exact details of the detection pipeline. We detect 2,301 aanns in the BabyLM corpus, which accounts for about 0.02% of all utterances, making construction quite rare for learners trained on this corpus.
Model
Our LMs are instances of the OPT LM (Zhang et al., 2022), which is an autoregressive transformer architecture. Our LMs have 12 layers and attention heads, use a vocabulary size of 16,384, and are trained for a maximum of 20 epochs using the transformers library (Wolf et al., 2020). The results we report for a given LM in our experiments are averaged over three different runs (with different random seeds). We list all other hyper-parameters and architectural details in Appendix A.
2.2 Acceptability data
To test our LMs on their knowledge of the aann, we use data from Mahowald (2023), which consists of 12,960 templatically generated sentences that contain aanns. Out of these, 3,420 contain acceptability ratings provided by 190 human participants, ranging from 1 (unacceptable) to 10 (acceptable). We use 7 as the threshold for clear acceptability, in that we only keep instances where human participants rated the acceptability of the construction in context to be greater than 7. We additionally discarded instances where the aanns appear in the BabyLM training set ($n=4$), as testing on these would not shed light on the LMs’ generalization behavior. This leaves us with 2,027 items.
For each aann instance in the dataset, Mahowald (2023) has also made available its corresponding corrupted versions, which focus on (1) adjective-numeral order; (2) presence of the article; (3) presence of the adjective; and (4) presence of the numeral. A hypothetical example of these corruptions is shown in Table 1 under the “aann” column. A model that has knowledge of the aann should find the well-formed instance to be more likely than each of its corrupted versions. Below we describe methods to measure likelihood and assess accuracy on these tests.
2.3 Measurements
Scoring
We use the Syntactic Log-odds Ratio (SLOR) proposed by Pauls and Klein (2012); Lau et al. (2017), to score sentences in our tests. Given a sentence containing a prefix followed by our target construction $\mathcal{C}{}$ and an optional suffix, slor is computed as the log of the ratio between the probability of the construction given the prefix as estimated by the LM, and that estimated by a unigram model, normalized by the length of the construction. Given a language model $m$ and a unigram estimator $u$:
$\textsc{slor}{}_{\text{prefix}}(\mathcal{C}{})=\frac{1}{\mid\mathcal{C}{}\mid}\log\frac{p_{m}(\mathcal{C}{}\mid\text{prefix})}{p_{u}(\mathcal{C}{})}$ (1)
Context aann anan naan
Well-formed a whopping ninety LMs a ninety whopping LMs ninety whopping a LMs
Corruptions
Order-swap a ninety whopping LMs a whopping ninety LMs whopping ninety a LMs
No article whopping ninety LMs ninety whopping LMs ninety whopping LMs
No modifier a ninety LMs a ninety LMs ninety a LMs
No numeral a whopping LMs a whopping LMs whopping a LMs
Table 1: Well-formed and corrupted examples of the aann construction and its counterfactual versions (anan and naan). Corruption types are taken from Mahowald (2023). Note that in the actual data, these constructions are embedded in full sentences.
Figure 2: Accuracies on tests for aann and its counterfactuals (anan and naan), achieved by LMs trained on BabyLM with various aann-manipulations (aann as is, no aann, anan, naan). The dashed line represents chance performance (6.25%), the dot-dashed line represents accuracies for slors computed using 2- and 4-gram LMs trained on BabyLM. Accuracies for GPT-2-XL (Radford et al., 2019) and Llama-2-7B (Touvron et al., 2023) are computed using log-probabilities, since unigram frequencies were unavailable for these LMs’ corpora.
Importantly, we train the unigram estimator for a given corpus using the same tokenizer used to train our autoregressive LMs on that corpus. We use slor in lieu of the usual normalized log-probability measure to ensure that the comparison between two models cannot be explained simply by the difference in unigram frequencies brought about by our manipulations. All log-probabilities were computed using minicons (Misra, 2022).
Accuracy
Using slor values computed over test instances, we calculate the accuracy of a model as:
$\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}\left[\bigcap\limits_{k}^{4}\left\{\frac{\textsc{slor}{}_{\text{prefix}_{i}}(\mathcal{C}{}_{i})}{\textsc{slor}{}_{\text{prefix}_{i}}(\mathcal{C}{}^{k}_{i})}>1\right\}\right],$ (2)
where, $\mathbbm{1}$ is the indicator function, and $k$ enumerates all four of our corruptions, and $N$ enumerates the test instances. That is, an instance within our test set is considered to be correct iff. the score on the well-formed construction, $\mathcal{C}_{i}$, is greater than the score for all corrupted instances, $\mathcal{C}^{k}_{i}$. The accuracy, then, is the proportion of correct instances within the test set. Since this involves four pairwise comparisons, chance performance is 6.25%.
2.4 Ablations
Common to subsequent experiments (Section 4 and Section 5) is the fact that we hold out certain parts of the BabyLM corpus—parts that conform to a certain linguistic or statistical hypothesis. This creates a gap between the experience of LMs trained on these ablated versions of the corpus, and that of the LM trained on the full BabyLM data. To circumvent this issue, we up-sample non-hypothesis-conforming utterances in BabyLM after performing our ablations, in a manner such that the LM still ends up encountering the exact same amount of tokens in such cases. As a result, all our LMs see the same quantity of tokens, but differ in content.
3 Experiment 1: LMs learn about aanns without having seen a single instance
3.1 LMs learn about aanns…
To investigate the extent to which LMs trained on BabyLM learn the aann construction, we measure their accuracy on our tests described in Section 2.2. From Figure 2, we observe that the BabyLM-trained LMs obtain accuracies around 70%, which is substantially above chance. This suggests that LMs can reasonably acquire generalizations to aanns from exposure to positive evidence that makes up only 0.02% of their training experience.
For comparison to larger, state-of-the-art LMs, we test Llama-2-7B (Touvron et al., 2023) and GPT-2 XL (Radford et al., 2019) on the aanns, and find them to achieve 83% and 78%, respectively. [^3] Qualitatively, performance is on these models is worse than data reported in Mahowald (2023), which used GPT-3. Because log-probabilities are not available from GPT-3, a direct comparison is not possible. Similarly, as a comparison to shallower LMs, we tested on 2-and 4-gram LMs trained on BabyLM and found both of them to achieve 0% accuracy, eliminating the possibility that the observed results are due to surface level statistics.
3.2 …without having seen a single instance…
Given that BabyLM-trained LMs learn the aann construction, how well would an LM generalize to aanns without having seen a single positive instance? To this end, we compare accuracies in the previous experiment to that obtained by LMs trained on BabyLM with our 2,301 detected aanns removed (i.e., no aann).
From Figure 2, we find LMs trained with the no aann condition to achieve an average accuracy of 54%, which is a noticeable drop compared to the 70% obtained by the LMs trained on the complete BabyLM corpus, but importantly 47.75 points above chance performance (and, as we show below, well above comparable baselines with other constructions). This is a non-trivial result, since it suggests that LMs are able to learn the acceptability of a construction’s instances, without having seen a single positive occurrence, which indicates that there exist some systematic patterns in the corpus that are driving this generalization behavior.
3.3 …more strongly than they learn counterfactual aann variants
To further contextualize the above results, we consider two counterfactual cases, where we replaced aanns in BabyLMs with instances involving the same lexical items, but in a word order that violates English grammar: (1) anan (e.g., a 90 whopping LMs); and (2) naan (e.g., 90 whopping a pages). This allows us to test if the results before are genuinely because LMs recognize the nuances of the aann construction, or whether they can learn alternative variants of the construction, which likely share fewer properties with other more frequent utterances in the data. If LMs are able to learn these constructions just as well as the the LMs in the previous experiments learned aanns, then the generalization claims from the previous experiments would be weakened.
To test for such possibilities, we create counterfactual versions of the Mahowald (2023) stimuli, where we apply analogous corruptions to the counterfactual variants of aann, such that they are violated in a similar manner as in the aann tests. Examples for the three types of instances in our tests can be found in Table 1. We then evaluate the previous two LMs (trained on BabyLM with and without seeing any aanns) with LMs trained on BabyLM with these counterfactual variants on judging the acceptability of aanns, anans, and naans. Figure 2 shows these results.
From Figure 2, we make two high-level observations. First, and most importantly, LMs that see anans and naans do not learn those constructions as well as they learn aanns—especially the LM that saw no aanns (54% aann accuracy compared to 36% naan accuracy obtained by the naan-trained LM). Second, these LMs end up learning aanns more strongly relative to how well they learn constructions that they observe in lieu of the aann—e.g., naan trained LM achieves around 54% accuracy on aanns even though naans appeared frequently in the data set and no aanns did. This, combined with the results of the previous two sub-experiments suggests strongly that LMs pick up on cues from other—likely related—constructions encountered during training in order to assign non-trivial probability to unseen instances of aanns.
Figure 3: slors on aanns from Mahowald (2023) for our LMs (left) and a 4-gram baseline (right) trained on BabyLM and its ablated versions. Our LMs show a range of hypothesized effects, whereas the 4-gram baselines only show an effect of the adjective/numeral balanced condition. slor for unablated BabyLM-trained LM shown with dotted line.
Phenomenon/Manipulation Example/Desc. Freq.
AANN a fine eighteen months 2,301
DT ANN
the usual forty dollars fine
14,347
A few/couple/dozen/etc. NNS
a few plums 55,226
Measure NNS with SG verbs
and/or indef. articles
6 months is a long time 62,597
A/An + ADJ/NUM balancing
enforce freq. balance
571,874
Random removal (control) randomized ablation 571,874
Table 2: Manipulated Phenomena, their examples/descriptions, and their frequency in the BabyLM corpus.
4 Experiment 2: Keys to Learning aanns
Our previous experiments reveal that, keeping everything else the same, LMs learn the aann construction far more accurately than they do its counterfactual variants. Furthermore, we also see strong aann acceptability judgments in LMs that have never encountered a single instance. This suggests that there are instances in the training data that contribute to the learning of the construction.
What might these be? Below we enumerate four hypotheses, each of which tackles subtly different aspects of the aann construction, then measure the effect of these phenomena, by separately holding them out during training and computing the average slor of the well-formed instances of the aann tests. The effect of a particular phenomenon on the acceptability of aanns can therefore be measured by comparing slors before and after holding out instances of that phenomenon. Methods for detecting the hypothesized phenomena can be found in Appendix B. As control, we additionally also hold out a random set of utterances, which do not conform to the target phenomena of interest. Note again that for all these cases, we ensure the LMs see the same number of tokens during training, by up-sampling other, non-aann containing sentences. Table 2 lists the hypotheses we consider, along with an example of their utterance and frequency of occurrence, in the BabyLM corpus.
The presence of “the ANN”
Phrases like “the beautiful five days” are common in corpora, which are not as unusual because “the” regularly takes plural nouns. We hypothesize that the acceptability of these structures affects the acceptability of aanns, since an LM might analogize from the general observation that ‘a’ or ‘an’ can substitute ‘the’ (e.g., a ball vs. the ball). Therefore, we consider all cases wheere a determiner precedes the contiguous sequence of article, numeral, plural noun.
A few/couple/dozen/etc. NNS
Another related phenomenon that is more common relative to the aann construction involves phrases such as “a few days” or “a couple bottles”. To an LM learner, they might provide evidence that measure noun phrases with plural nouns can be attached to an indefinite article (a/an), as is the case in aanns.
Measure NNS treated as singular
We consider yet another phenomenon involving phrases that treat measure nouns as singular, this time in terms of agreement, e.g., “Five miles is a long way to go”, and “1,000 pages is a lot for a dissertation.” These cases might provide further evidence to the model that measure noun phrases with plural nouns can be treated as a singular unit, thereby affecting the acceptability of the aann. These are less prevalent compared to the cases involving a few/couple/dozen NNS, but still far more frequent than the aann, therefore, we combine the two as a general case of treating measure NPs as singular.
Balancing the frequencies of A/An + ADJ/NUM
A more surface-level reason why “a beautiful five days” might be more natural to LMs than is “a five beautiful days”, could be that adjectives are more likely to follow indefinite articles than are numerals. For instance, adjectives are $\approx$14.6 times more likely to follow indefinite articles in the BabyLM corpus than are numerals. To measure this effect, we hold out instances such that adjectives and numerals are equally likely to follow an indefinite article. This ends up being the largest portion of the data that we hold out.
Control: Random removal
A potential confound in the above ablations could be that the slor values of the aanns go down merely due to loss of content—this could be despite the fact that we add back additional tokens from BabyLM (such that all LMs see the exact same amount of tokens). To account for this, we additionally consider a control condition where we remove as many tokens as in the largest ablation (i.e., the A/An + ADJ/NUM case) such that none of the above phenomena are taken out. Insofar as the LMs rely on the aforementioned phenomena, results on LMs trained with this ablation should be closer to the original BabyLM-trained LM’s results.
4.1 Analysis and Results
In our experiments, we individually ablate out each of the aforementioned phenomena under two settings: (1) aanns are removed during training in addition to the phenomena; and when possible, (2) aanns are seen during training. (1) is a stricter setting, since here the LMs see neither the target phenomenon nor a single instance of the aann. Comparing average slors obtained in this condition to that obtained for the no aann can shed light on the the extent to which the target phenomenon is critical in allowing LMs to assign non-trivial probabilities on unseen aanns, zero-shot. On the other hand, (2) still allows for the LMs to perform lexical generalization (Kim et al., 2022) from seen aann instances—where they may arrive at strong acceptability measures on the test aanns by performing slot filling, without necessarily relying on the hypothesized phenomena (characterization of which is equally important).
Figure 3 shows the average slors obtained across various ablations under the two settings. As a baseline, we compare our results to 4-gram LMs [^4] Trained using KenLM (Heafield, 2011) trained on corresponding ablations of the BabyLM corpus. We observe that holding out all our hypothesized phenomena has non-trivial effects on our LMs’ ratings of unseen well-formed aanns, and that these effects are different for when aanns are seen during training, or are held out. When aanns are held out along with the phenomena, we see substantial drops in the average slor values assigned by LMs on unseen aanns relative to that assigned by LMs in the no aann condition. Specifically, balancing the frequency of adjectives and numerals following an article has the greatest effect, followed by the two cases where measure nouns are treated as singular, followed by the removal of all cases that involve determiner + adjective + numeral + noun sequences. This suggests that in the absence of even a single aann during training, these phenomena are critical for LMs to assign non-trivial likelihoods to aanns. Importantly, simply ablating large amounts of data cannot explain these results, since LMs trained on our controlled condition show higher slor values than in all our hypothesis-informed ablations. These patterns are absent in 4-gram LMs, suggesting that they do not arise as a result of shallow, surface statistics, with the slight exception of differences observed in the case of the article+adjective/numeral ablation. Overall, this is a noteworthy finding, since it indicates that LMs can demonstrate a completely novel phenomenon (aann) by relying on other related—and more frequent—phenomena.
When aanns are seen during training, however, we observe LMs’ results on unseen aanns to show more similar slor values with respect to the the LMs trained on the unablated BabyLM corpus, although they are still significantly reduced in some cases (e.g., singular measure nouns).
We conclude that when LMs see evidence of the aann construction, they do learn from it. But key related phenomena where measure nouns are treated as singular do seem to show some notable effects even when AANNs are present, suggesting that these enable additional learning even when aanns are present.
5 Experiment 3: The Role of Variability
Results from the previous experiment highlight the importance of seen aanns in order for LMs to generalize to unseen instances. What properties of these seen instances make LMs generalize? More broadly, there is a longstanding question as to how the nature of the instances of a construction provided during learning affect its (partial) productivity (Goldberg, 2005, 2019). In the context of aanns, we consider the role of variability on the open slots of the construction as a factor that might play a role in LMs’ productivity on unseen instances. The idea that slot-variability could affect learnability is motivated by theoretical claims in usage-based linguistics (Bybee, 1995), as well as existing research on novel constructions (Suttle and Goldberg, 2011), morphological productivity (Baayen, 2009; O’Donnell, 2015), and inductive inferences in cognitive psychology (Osherson et al., 1990; Xu and Tenenbaum, 2007). The idea is that encountering a slot with a wide variety of lexical items serves as a cue that the slot is flexible.
We hypothesize that instances of aanns that provide natural evidence of greater open-slot variability—i.e. evidence that many different adjectives, numerals, and nouns can go in their respective positions in the aann construction—would lead LMs to assign greater likelihood to unseen aanns. On the other hand, LMs that only encounter a restricted set of instances might overfit, and be more conservative in extending the coverage of possible aanns to novel combinations of the slot-fillers. To test this, we divided our set of 2,301 aann-containing utterances in the BabyLM corpus into two equal subsets—one that contained aanns which were individually highly frequent but restricted in the types of adjective/numeral/nouns, and the other where the aanns were individually less frequent, but showed more variability in those slots. We then trained LMs on the BabyLM corpus containing utterances involving either of these two cases. Figure 4 shows the resulting average slors obtained from this experiment, along with those obtained by LMs trained on unablated BabyLM and the no aann conditions.
From Figure 4, we see that LMs that only saw aanns that were highly variable in their open-slots demonstrated slors that were comparable and sometimes greater than those obtained by LMs that saw all aanns. By contrast, LMs that only saw aanns with low variability were as good as LMs that never saw any aanns. Therefore, LMs were sensitive to the nature of range of fillers that went into the construction’s open slots, showing relatively greater productivity when they observed evidence that the slots were highly variable. This is compatible with our hypothesis that slot-variability might affect the extent to which LMs “permit” productive uses of a construction.
Figure 4: slors on aanns from Mahowald (2023) for LMs trained on BabyLM with low and high variability in the observed instances of aann. slor for unablated BabyLM-trained LM shown with dotted line.
6 Conclusion
There is increasing interest in computational linguistics in how language models can handle what has been variously called the “long tail” of language (Prange et al., 2021), “extremely rare constructions” (Potts, 2023), “exceptions to syntactic rules” (Leong and Linzen, 2023), “rare linguistic phenomena” (Weissweiler et al., 2024), inter alia. Studies of such phenomena are important first because LLMs (and LMs and statistical models in general) are known to be extremely sensitive to frequency and to perform far better in data-rich environments and, second, because the human ability to generalize to rare phenomena is central to knowledge of language.
We found that LMs trained on a human-scale of data can learn a rare construction like the aann. We found that this learning occurs even without veridical instances of the aann construction in the training data, and that it is mediated by occurrences of other related constructions in training. As such, these results join a growing body of work showing the ability of LLMs to learn constructions (Tayyar Madabushi et al., 2020; Tseng et al., 2022; Li et al., 2022; Veenboer and Bloem, 2023).
7 Limitations
In future work, it would be valuable to extend this method to a wider range of constructions. But scaling this approach up is not straightforward since it requires identifying and extracting idiosyncratic constructions, and—more onerously—developing testable hypotheses about what makes them learnable from limited amounts of data. While this is a limitation, it also calls for more synergistic collaborations between theoretical and computational linguists.
Another limitation is that our method requires repeated training of LMs from scratch which can be computationally expensive. Alternate methods could be to ablate knowledge of particular hypotheses using representational editing methods such as AlterRep (Ravfogel et al., 2021), etc.
Unlike Weissweiler et al. (2022), we do not test the ability to interpret these constructions for downstream tasks. Instead, we focus on linguistic form. Extending these results to semantic tasks would be informative.
8 Acknowledgments
(KM)2 acknowledge funding from NSF Grant 2104995 awarded to Kyle Mahowald. For helpful conversations, we thank Adele Goldberg, Leonie Weissweiler, the computational linguistics research group at UT Austin, the syntax-semantics research group at UT Austin, the audience at the Texas Linguistics Society meeting. We thank Chris Potts for his paper on the PiPP construction (Potts, 2023) which inspired the “keys to all of this” idea.
Appendix A LM training details
As mentioned in the main text (see Section 2), we use the OPT architecture (Zhang et al., 2022) to train our LMs on all versions of the BabyLM corpus. This was the best performing autoregressive LM in the BabyLM Competition (Warstadt et al., 2023). For each instance of the BabyLM (ablated or otherwise), we tune the learning rate [^5] We searched the following set: {1e-4, 3e-4, 1e-3, 3e-3} based on the validation set, and use the best learning rate as a result of the tuning to train an additional two language models using different seeds. As a result, for each ablation of the BabyLM corpus, we run 6 LM training experiments, which amounts to a whopping 90 LMs for all our experiments. Table 3 contains further details of the training.
(Hyper)parameter Value
Architecture OPT (Zhang et al., 2022)
Embed size 768
FFN dimension 3,072
Num. layers 12
Attention heads 12
Vocab size 16,384
Max. seq. length 256
Batch size 32
Warmup steps 32,000
Epochs 20
Total parameters 97M
Training size 100M tokens
Compute 1x NVIDIA A40
Training time 21 hours
Table 3: LM Training details
Appendix B Detecting aanns and related phenomena
In this section, we briefly describe our methods to extract constructions and phenomena relevant to this paper from the BabyLM corpus (Warstadt et al., 2023). Our methods primarily rely on: 1) the surface form of the sentences in the corpus; 2) their corresponding part-of-speech (POS) tag sequences; and in a few cases, 3) their depedency parses. For the latter two, we used spacy (Honnibal et al., 2020), specifically, its en_web_trf model, which is based on the RoBERTa-base LM (Liu et al., 2019). Next we describe how we used these artifacts to detect our target constructions:
B.1 AANNs
To detect AANNs we primarily rely on POS-tagged sequences, and construct a regex pattern over them [^6] In reality this was constructed over several iterations, taking into account many different possible realizations of the construction in free text. which is able to robustly detect aanns:
[…]
where we restrict the determiner (DT) to be either ‘a’, ‘an’, or ‘another’. This regex permits multiple adjectives (an exhilarating and marvelous three months) optional adverbs (an excruciatingly painful two semesters), multi-word noun phrases with plural head-nouns (a refreshing two glasses of aperol spritz), numeral-expressions involving subordinate clauses (a measly three to five days), among other potential edge cases. We additionally use the following adjectives as proxies for numerals, as per the guidelines of Kayne (2007) and Solt (2007):
numeral_proxies = [’few’, ’dozen’, ’couple’, ’several’, ’many’, ’more’]
For instance, the following examples would all count as instances of the aann:
[…]
Once detected, we map the found constructions to their respective positions within the aann format, which allows us to measure metrics such as slot variability, etc.
B.2 DT ANNs
We follow the exact same procedure as the one for aanns, but no longer restrict the determiner position to only be an indefinite determiner.
B.3 A few/couple/dozen NOUNs
An important phenomenon that we consider to be related to the aann involves cases such as: “that only lasted a few days” and “could you bring me a couple liters?”, etc., where the plural nouns are attached to an indefinite article. To detect such cases, we consider the following two dependency configurations, where we have an indefinite determiner (a, an, another) with either a det relation with the plural noun (NNS or NNPS) or a quantmod relation with a noun which has a nummod with the plural noun. In the former case, we usually have an amod relation between the noun and the adjective.
{dependency}{deptext}
[column sep=0.6cm, row sep=.1ex] …& DT & JJ & NNS & …
…& a & few & days & …
[…]
{dependency}{deptext}
[column sep=0.6cm, row sep=.1ex] …& DT & NN & NNS & …
…& a & couple & days & …
[…]
B.4 Measure NNS with Singular Verbs
Similar to the previous case, another phenomenon which might be related to the aann constructions is when measure noun-phrases with plural nouns are treated as singular via their agreement with a verb—e.g., “five dollars is plenty!” To detect such cases, we again rely on the following dependency configuration, where we have a plural noun (NNS or NNPS) attached to a cardinal number (CD) via the nummod dependency relation, and at the same time also attached to singular verbs via the nsubj dependency relation (i.e., are subjects of singular verbs).
{dependency}{deptext}
[column sep=0.6cm] …& CD & NNS & VB & …
…& five & dollars & is & …
[…]
Appendix C A/An + ADJ/NUM frequency balancing
A corpus analysis of BabyLM, along with its POS-tagged version suggests that the sequence “a/an/another (JJ|JJR|JJS)” occurs 613,985 times while “a/an/another CD” occurs only 42,111 times – this suggests that adjectives are approximately 14.6 more likely to follow an indefinite article than are numerals. We therefore balance these values by removing 571,874 instances where adjectives follow an indefinite article. This constitutes the largest-sized ablation in this work.